# Funding AI for Good: A Call for Meaningful Engagement

### v1

@@ 2.2. Lessons from Prior “Tech for Good” Projects: Techno-Centric vs Balanced Approaches
In this section, we synthesize this rich body of work and categorize characteristics of projects into a more techno-centric approach (bias towards technical solutions) and a more balanced approach (emphasizing both technical capacities and contextual understanding). We particularly focus on aspects of motivations, collaborations, longevity and process, and evaluations, because these aspects are closely influenced by funding agendas (Lin et al., 2024). While prior work has pointed out a techno-centric orientation of some AI4SG initiatives, this paper offers a granular analysis of how this orientation is instantiated across specific components of funding documents.
@@ 2.2.1. Techno-Centric Approaches
In “tech for good”, techno-centric projects are driven by an “innovate at all costs” mentality, leading to the creation of technologies with little maintenance support and local capacity building afterward (Saha et al., 2022b). It incentivizes a top-down design approach, where the solutions are typically designed by external technologists without adequate engagement with the local communities (Walsham, 2017; Dada, 2006; Harris, 2016; Dodson et al., 2012), especially during the decision-making stages of projects (Saha et al., 2022a). It could also lead to the phenomenon of “bungee jumping research” (Dearden and Tucker, 2016), characterized by brief, superficial engagement (Saha et al., 2022b; Sultana et al., 2019; Dada, 2006). Development of the solutions tends to follow a linear and short-term process: building the technology, deploying it, and moving on to the next project, where researchers often leave after their work is published, making it difficult to sustain the systems they have introduced (Taylor et al., 2013; Dada, 2006; Walsham, 2017). Last but not least, evaluations of techno-centric projects constrain assessments of success to technical dimensions, narrowly focusing metrics such as model accuracy or the production of tangible artifacts like software (Dodson et al., 2012; Lin et al., 2024).
@@ 2.2.2. Balanced Approaches
2.3. Gaps in Our Understanding of AI4SG Funders and Funding Agendas
@@
Despite the central role that funders play in shaping the AI4SG ecosystem, there is limited scholarship examining how funders frame AI4SG and what approaches are envisioned by funders. An emerging body of work has started to understand the perspectives, incentives, and challenges of grantmakers in the broader “tech for good” space. For example, through interviews with eight decision-makers in development projects, Saha et al. (2022a) found that community participation is rarely a part of the grant decision-making process, and it is not a requirement for applicants either. Some barriers to incorporating community voices include limited access to local organizations and communities and overconfidence in their knowledge about the local contexts (Saha et al., 2022a). Relatedly, Gardner et al. (2022) point out that a lack of internal expertise in AI and AI ethics creates barriers for funders to incorporate Trustworthy AI principles. Studies on other influential decision-makers, such as policymakers in government agencies, point out factors influencing their decision-making process around innovative technologies, including learning from neighboring agencies’ experiences, reliance on national surveys (which omit marginalized communities’ views of their needs) (Saha et al., 2022b). These decision-makers also face obstacles in effective community engagement, including power and knowledge differentials and inadequate infrastructural support (Kawakami et al., 2024a).
@@
Focusing on private tech companies specifically, some recent works have started to critically examine investment in AI4SG (Nost and Colven, 2022; Lukacz, 2024; Henriksen and Richey, 2022). For example, Nost and Colven (2022) view Microsoft’s AI for Earth program as a form of greenwashing and “philanthrocapitalism”, where the company uses the initiative to enhance its reputation as a socially responsible corporation, while simultaneously contributing to environmental harm through its broader business operations, including its reliance on fossil fuels for data centers. Lukacz (2024) further argues that Microsoft’s narrative of democratizing AI and environmental data obscures the power they wield in shaping the technology and its impacts. Henriksen and Richey (2022) examine Google’s philanthropic efforts, arguing that they serve dual purposes: advancing social causes while simultaneously expanding Google’s market influence and integrating its technologies into more areas of life. Comparing “Big Tech” companies with the “Big Tobacco” industry, Abdalla and Abdalla (2021) examined “Big Tech” funding’s influence on the broader academic landscape (though not specific to AI4SG), impacting events and research questions that “put forward a socially responsible public image” of tech companies.
@@ 3. Methods
To examine how funders conceptualize AI4SG and what approaches they envision in AI4SG, we analyzed publicly available funding documents. In this section, we detail our approach for identifying and analyzing AI4SG funding documents (Subsections 3.1 and 3.2) as well as our team’s positionality (Subsections 3.3). We discuss limitations to our approach and opportunities for future work in Subsection 5.4.
@@ 3.1. Qualitative Dataset Development
In this paper, we define “funders” as organizations that provide monetary support to AI4SG, such as philanthropic foundations, nonprofit organizations, government agencies, and private technology companies. While individuals may also fund AI4SG projects, we focused on organizations because they are more likely to provide public articulations of their funding agendas. We define “funding programs” as specific AI4SG initiatives, such as a research grant on the intersection of climate change and AI. We define “funding documents” as funding calls for proposals and grant announcements through which funders communicate their intentions and envisioned approaches in AI4SG. While funding calls typically outline more detailed guidelines, grant announcements present a more concise overview of an initiative’s aims, scope, and thematic priorities. We consider both as funding documents because they offer complementary insights into funders’ intentions.
@@
We used the Google search engine to identify publicly available AI4SG funding documents published prior to August 2025. Following the approach in Wong et al. (2023) when searching for AI ethics toolkits, we sought to emulate how AI4SG stakeholders (e.g., research teams, community-based organizations, individual practitioners) might seek funding opportunities. This approach also led us to discover and examine webpages from grant databases, such as nsf.gov/funding/opportunities.
@@
  • Funding documents must specify an orientation towards developing, using, or evaluating AI, Machine Learning, Data Science, or other data-driven technologies for social applications. This excluded projects on AI education (e.g. (NSF, 2023)), which focus on educating people about AI instead of exploring AI as an application in the education field (e.g., AI assistants in classrooms).
@@
3.2. Qualitative Analysis
We conducted a thematic analysis (Braun and Clarke, 2006, 2019) of the funding documents through an iterative coding process. In the first stage of coding, the first two authors coded basic information about the funding programs, including the methods, the social issues, the duration of projects, the intended applicants, and the amount of funding. In the next coding stage, the first two authors independently conducted an open-ended inductive qualitative analysis of three randomly chosen funding documents and discussed the coding strategies. Through iterative discussions, the two authors identified some initial themes, which helped refine our research questions and narrow the scope of analysis to four main components of funding documents: 1) Background, 2) Expected Outcomes, 3) Eligibility Criteria and Team Composition, and 4) Funders’ Support. We also kept detailed memos as we coded the themes to document our observations. In the third coding stage, the first author coded all the funding documents, with our research questions in mind. Each funding document’s codes were second reviewed by another author in our team. The first author then resolved any discrepancies with the rest of the team through online discussions. All authors then conducted iterative affinity diagramming on an online whiteboard and discussed emerging high-level themes from the codes. We collaboratively characterized these high-level themes into a spectrum between a more techno-centric approach or a more balanced approach using the lens of prior literature synthesized in Section 2.2. Appendix B provides screenshots of our coding process and affinity diagrams.
@@ 4. Findings
In what follows, we report on themes in different components of funding documents and highlight rhetoric leaning towards a more techno-centric or balanced approach to AI4SG. We begin with the Background (Section 4.1), where we discuss how funders frame “social good”, “AI”, and how AI is envisioned to help address social good problems (RQ1). In the Expected Outcomes and Evaluations (Section 4.2), we discuss funders’ envisioned outcomes, often communicated through evaluation criteria (RQ2). In the Eligibility (Section 4.3) and Support from Funders (Section 4.4) sections, we shed light on how funders envision “AI4SG” can be achieved and by whom (RQ3). Table LABEL:tab:funding_doc_mr in Appendix A provides a detailed overview of our dataset, including funding amount, duration, targeted regions, funder types, and target regions.
@@ 4.4.2. More balanced
Tailoring support activities to the needs of applicants. Some other funding programs provided broader forms of support, beyond just technical resources and training. For example, as part of the Wellcome Mental Health Data Prize (C8 and C9), Wellcome Trust offered a web-based platform to facilitate knowledge-sharing across teams. The funder also facilitated matchmaking events where “individuals with different knowledge, skills and backgrounds will be introduced to collaborate and develop a proposal together” to enable interdisciplinary collaborations. In another funding call (C7), the Mozilla Foundation described a broad array of support that funded teams would receive: “While the specific types of support are tailored to the unique needs of the awardees and cohorts, providing opportunities for peer learning and community, mentorship, and communications support are key elements of our accompaniment approach.” However, none of the funding documents in our corpus explicitly mentioned training resources and support for local community engagement and participatory design.
@@ 5. Discussion
Our analysis presents a range of framing and approaches envisioned by AI4SG funders through funding documents. In this section, we first highlight the disconnects between the positive social impact that AI4SG aims for and the techno-centric approaches that some funding documents promoted (Section 5.1). Building on the more balanced approaches and existing gaps identified in our dataset, we then propose additional opportunities for designing future funding calls, including supporting relationship building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, emphasizing the role that the HCI community can and should play in helping shape AI4SG funding agendas (Section 5.3). We discuss the limitations of our work and lay out several future research directions in Section 5.4.
@@
5.1. Revealing Dissonances Between Intended Outcomes and More Techno-centric Approaches
@@
Many AI4SG funding programs in our dataset expected feasible and sustainable deployment of AI technologies or novel research advancement in understanding AI’s potential (Section 4.2). However, more techno-centric approaches promoted through some funding documents may lead to outcomes that fall short of AI4SG’s aim for positive and long-lasting change in social domains (Perera, 2024; Tudor et al., 2024). We highlight three prominent aspects below.
@@
First of all, funding documents that motivated AI’s use with large-scale social problems without instantiating concrete contextual benefits (Section 4.1.1) could lead to premature emphasis on scalability, risking putting such a strong emphasis on creating something that works everywhere that it ends up not working particularly well anywhere. As Hanna and Park note, “scale thinking” necessitates homogenizing users into predefined categories, disregarding the diversity of local contexts (Hanna and Park, 2020). The urgency to scale and short-term funding could incentivize rushed implementations that are poorly suited to diverse contexts (Heeks, 2003, 2002; Dada, 2006). As Burrell and Toyama (2009) put it, “in an effort to solve problems quickly on large scale, a concern for local values, preferences, and opinions may become secondary,” cautioning against favoring only large-scale, generalizable studies at the expense of in-depth, context-rich work (and vice versa). We do not argue that scalable and generalizable approaches should not be pursued ever, but we suggest that they should be pursued inductively. That is, AI4SG initiatives should focus early efforts on solving specific problems of specific communities effectively and sustainably, and then look for commonalities that could inform the development of generalizable technical capabilities and infrastructures.
@@
Secondly, funding rhetoric that led with assumptions of AI’s deterministic benefits without acknowledging its risks (Section 4.1.1) could reinforce the wider hype and “enchanted determinism” surrounding AI (Campolo and Crawford, 2020). This narrative fosters unrealistic expectations (Lin et al., 2024; Taylor et al., 2013), promotes techno-solutionism, and diverts attention from underlying social issues (Walsham, 2017; Karusala and Kumar, 2017). Furthermore, when funding documents framed AI as a force of good for the Global South or nonprofit sectors deterministically (Section 4.1.1), it risks promoting top-down approaches that cast these groups as lacking and in need of catching up (Erete et al., 2023; Valencia, 1997; Dearden and Tucker, 2016). This framing ignores the unique needs (and assets) of public sectors and the Global South (Hanna and Park, 2020). While corporate donations of computing resources and support (Section 4.4.1) can expand access, they may also reinforce Big Tech’s hegemonic narrative (Collins, 2000; Lukacz, 2024; Nost and Colven, 2022; Seger et al., 2023), addressing surface-level gaps rather than deeper structural issues (Ho et al., 2009b; Burrell and Toyama, 2009; Dearden and Tucker, 2016). These initiatives risk reproducing patterns of saviorism, positioning the Global South as a testing ground for foreign-owned systems while extracting valuable data (Damilare Dosunmu, 2025). These projects also overlook the uneven geography of the AI supply chain; the Global South is already part of AI, but primarily through the underpaid labor that makes AI systems work (Posada, 2022).

### v2

@@ 2.2. Lessons from Prior “Tech for Good” Projects: Techno-Centric vs Balanced Approaches
In this section, we synthesize this rich body of work and categorize characteristics of projects into a more techno-centric approach (bias towards technical solutions) and a more balanced approach (emphasizing both technical capacities and contextual understanding). We particularly focus on aspects of motivations, collaborations, longevity and process, and evaluations, because these aspects are closely influenced by funding agendas (Lin et al., 2024).
@@ 2.2.1. Techno-Centric Approaches
In “tech for good”, techno-centric projects are driven by an “innovate at all costs” mentality, leading to the creation of technologies with little maintenance support or local capacity building afterward (Saha et al., 2022b). It incentivizes a top-down design approach, where the solutions are typically designed by external technologists without adequate engagement with the local communities (Walsham, 2017; Dada, 2006; Harris, 2016; Dodson et al., 2012), especially during the decision-making stages of projects (Saha et al., 2022a). It could also lead to the phenomenon of “bungee jumping research” (Dearden and Tucker, 2016), characterized by brief, superficial engagement (Saha et al., 2022b; Sultana et al., 2019; Dada, 2006). Development of the solutions tends to follow a linear and short-term process: building the technology, deploying it, and moving on to the next project, where researchers often leave after their work is published, making it difficult to sustain the systems they have introduced (Taylor et al., 2013; Dada, 2006; Walsham, 2017). Last but not least, evaluations of techno-centric projects constrain assessments of success to technical dimensions, narrowly focusing on metrics such as model accuracy or the production of tangible artifacts like software (Dodson et al., 2012; Lin et al., 2024).
@@ 2.2.2. Balanced Approaches
2.3. Emerging Understanding of “Tech for Good” Decision-Makers and AI4SG Investments
@@
An emerging body of work has begun to understand the perspectives, incentives, and challenges faced by grantmakers in “tech for good.” For example, through interviews with eight decision-makers in development projects, Saha et al. (2022a) found that community participation is rarely a part of the grant decision-making process, and it is not a requirement for applicants either. Some barriers to incorporating community voices include limited access to local organizations and communities, and overconfidence in their knowledge about the local contexts (Saha et al., 2022a). Relatedly, Gardner et al. (2022) argue that a lack of internal expertise in AI and AI ethics creates barriers for funders to incorporate Trustworthy AI principles. Studies on other influential decision-makers, such as policymakers in government agencies, point out factors influencing their decision-making process around innovative technologies, including learning from neighboring agencies’ experiences and reliance on national surveys (which omit marginalized communities’ views of their needs) (Saha et al., 2022b).
@@
Focusing on private tech companies specifically, some recent works have started to critically examine investment in AI4SG through case study analysis (Nost and Colven, 2022; Lukacz, 2024; Henriksen and Richey, 2022). For example, Nost and Colven (2022) view Microsoft’s AI for Earth program as a form of greenwashing and “philanthrocapitalism”, where the company uses the initiative to enhance its reputation as a socially responsible corporation, while simultaneously contributing to environmental harm through its broader business operations, including its reliance on fossil fuels for data centers. Henriksen and Richey (2022) examine Google’s philanthropic efforts, similarly arguing that they serve dual purposes: advancing social causes while simultaneously expanding Google’s market influence and integrating its technologies into more areas of life. Comparing “Big Tech” companies with the “Big Tobacco” industry, Abdalla and Abdalla (2021) examined “Big Tech” funding’s influence on the broader academic landscape (though not specific to AI4SG), shaping events and research questions that “put forward a socially responsible public image” of tech companies.
2.4. Document Analysis of Pre-Existing Texts about AI and Funding (beyond AI)
Document analysis is a systematic procedure for evaluating printed or electronic documents to “elicit meaning, gain understanding, and develop empirical knowledge” (Bowen, 2009). Documents are pre-existing textual instruments (e.g., public reports, proposals, meeting notes) that serve as “social facts” (Bowen, 2009) and can be qualitatively evaluated (Morgan, 2022). Document analysis can provide “contextual richness” by offering background information and inspiring new research questions (Bowen, 2009). In addition, because documents are “unaffected” by researchers’ presence and influence (unlike interview data), document analysis provides a way to evaluate stable texts that are not distorted by participant reactivity and thus offer access to information that may be difficult or impossible to gather through interactive methods (Morgan, 2022). Although pre-existing texts only contain “limited information” and reflect only what their creators (such as funders) choose to make available, they still provide a meaningful part of the larger picture and guide future research directions (Morgan, 2022).
HCI scholars have used document analysis to examine pre-existing texts about AI, encompassing AI ethics toolkits (Wong et al., 2023b), AI policy documents (Wong et al., 2023a; Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), AI risk manifesto-style statements (Oldenburg and Papyshev, 2025), patents (Cheon, 2023), and military funding (Widder et al., 2024). Pre-existing documents about AI do not simply describe AI as facts; they construct AI imaginaries—“collective visions, beliefs, symbols, and expectations that individuals and communities hold about AI and its potential outcomes” (Zhong et al., 2025), rooted in Jasanoff and Kim’s “sociotechnical imaginaries” (collective visions of technology’s role in society) (Jasanoff and Kim, 2009). These documents not only define what AI is, what problems it should solve, who benefits from it, and whose perspectives remain invisible, they also “actively shape the trajectory of AI development, integration, and governance” (Zhong et al., 2025). For example, Wong et al. (2023b) analyzed a corpus of 27 AI ethics toolkits to “identify the discourses about ethics, the imagined users of the toolkits, and the work practices the toolkits envision and support.” They found that AI is imagined as a technology that can be made “ethical” through technical adjustments—better models, checklists, or audits (Wong et al., 2023b)Sophie Bennani-Taylor (2024) argue that policy documents perform “discursive infrastructuring”: they stabilize certain meanings of AI and legitimize particular sociotechnical arrangements. Through an analysis of UK’s National AI Strategy document, the author found that AI is constructed as 1) an autonomous, inevitable force, 2) a driver of national progress and economic modernization, and 3) a domain dominated by technical expertise over social knowledge (Sophie Bennani-Taylor, 2024). Closest to the corpus analyzed in our work, Widder et al. (2024) analyzed U.S. Department of Defense’s grant solicitations for the use of AI in military applications and revealed a deep entanglement between academic AI research and militarized imaginaries, where recurring rhetorical devices like “one small problem” justify continuous funding increases for military AI.
Similarly, prior research examining funding documents (not limited to AI) shows that funding languages are not neutral; they encode priorities, shape applicant behaviors, and structure downstream research collaborations. For example, Palmero et al. (2025) conducted a textual network analysis of Horizon Europe funding calls, treating these texts as “cognitive artefacts of policy design” that reveal how high-level policies are translated into operational agendas. The authors point out that funding document “predefines applicants’ behaviour, shapes their own perceptions of eligibility, and influences evaluative processes” (Palmero et al., 2025). Existing work in innovation studies likewise demonstrates the downdream influence of funding documents: Peng et al. (2024) find that the proportion of ”promotional language” (such as “innovative” and “unique”) in grant proposals is statistically associated with a proposal’s likelihood of being funded, while Smith et al. (2023) show “how thematic directives rooted in funding calls influenced collaboration across research.”
Building on this body of work, we contribute a first document analysis of AI4SG funding documents to understand how funding language discursively constructs AI imaginaries (especially AI’s role in social good domains) and characterizes the conditions under which its positive impact might be achieved.
@@ 3. Methods
To examine how funding documents frame AI4SG, we collected and analyzed funding calls and grant announcements that are publicly accessible to AI4SG project applicants and the wider public audience. In this section, we detail our approach for identifying and analyzing AI4SG funding documents (Subsections 3.1 and 3.2) as well as our team’s positionality (Subsections 3.3). We discuss limitations to our approach in Subsection 3.4.
@@ 3.1. Qualitative Dataset Development
In this paper, we define “funders” as organizations that provide monetary support to AI4SG, such as philanthropic foundations, nonprofit organizations, government agencies, and private technology companies. While individuals may also fund AI4SG projects, we focused on organizations because they are more likely to provide public articulations of their funding agendas. We define “funding programs” as specific AI4SG initiatives, such as a research grant on the intersection of climate change and AI. We define “funding documents” as funding calls for proposals and grant announcements through which funders communicate their intentions and envisioned approaches in AI4SG. While funding calls typically outline more detailed guidelines, grant announcements present a more concise overview of an initiative’s aims, scope, and thematic priorities. We consider both as funding documents because they offer complementary insights into funding agendas.
@@
We used the Google search engine to identify publicly available AI4SG funding documents published prior to August 2025. Following the approach in Wong et al. (2023b) when searching for AI ethics toolkits, we sought to emulate how AI4SG stakeholders (e.g., research teams, community-based organizations, individual practitioners) might seek funding opportunities. This approach also led us to discover and examine webpages from grant databases, such as nsf.gov/funding/opportunities.
@@
  • Funding documents must mention developing, using, or evaluating AI, Machine Learning, Data Science, or other data-driven technologies for social applications—ranging from AI as a tool to be assessed or scrutinized to AI as a proposed solution. This excluded projects on AI education (e.g. (NSF, 2023)), which focus on educating people about AI instead of exploring AI as an application in the education field (e.g., AI assistants in classrooms).
@@
3.2. Thematic Analysis on Funding Documents
We conducted a reflexive thematic analysis (Braun and Clarke, 2006, 2019) of the funding documents through an iterative coding process. In the first stage of coding, the first two authors coded basic information about the funding programs, including the methods, the social issues, the duration of projects, the intended applicants, and the amount of funding.
In the next coding stage, the first two authors independently conducted an open-ended inductive analysis of three randomly chosen funding documents and discussed the coding strategies. Through iterative discussions, the two authors identified some initial themes, which helped refine our research questions and narrow the scope of analysis to four main components of funding documents: 1) Background, 2) Expected Outcomes, 3) Eligibility Criteria and Team Composition, and 4) Funders’ Support. We did not produce a fixed codebook because reflexive thematic analysis treats coding as an iterative, flexible, and interpretive process rather than a procedure of reliably categorizing data into predefined buckets (Braun and Clarke, 2019). Instead, we kept detailed memos as we allowed the themes to develop inductively through ongoing, reflexive engagement with the data, researchers’ memos, and discussions among the research teams.
In the third coding stage, the first author coded all the funding documents, with our research questions in mind. Each funding document’s codes were reviewed by another author in our team. The first author then resolved any (though few) disagreements (denoted as comments on Google documents) with each code reviewer through online discussions. These discussions involved clarifying the meaning of the first author’s original codes and discussing whether they aligned with the original texts in the funding documents. We did not calculate intercoder reliability, since reflexive thematic analysis treats coding as an interpretive process rather than one that seeks positivist (quantitative) reliability measures (Braun and Clarke, 2019). All authors then conducted iterative affinity diagramming on an online whiteboard and discussed emerging high-level themes from the codes. The purpose of affinity diagramming was to categorize low-level codes (in the sentence level) and surface higher-level themes across the funding documents (in the section level and beyond), as well as finding connections between these themes. We collaboratively characterized these high-level themes into a spectrum between a more techno-centric approach or a more balanced approach using the lens of prior literature synthesized in Section 2.2. Appendix B provides screenshots of our coding process and affinity diagrams.
@@ 3.3. Positionality
3.4. Limitations
The list of funding documents in our dataset is neither exhaustive nor statistically representative. Our analysis is limited to English-language materials, which excludes funding programs described in languages other than English. Additionally, funding programs that do not use the keywords in our study might have been overlooked. Finally, we emphasize that the funding documents in our corpus fall along a spectrum from more techno-centric to more balanced approaches, and we caution readers against viewing this spectrum as a dichotomy.
@@ 4. Findings
In what follows, we report on themes in different components of funding documents and highlight rhetoric leaning towards a more techno-centric or balanced approach to AI4SG. We begin with the Background (Section 4.1), where we discuss how funding documents framed “social good”, “AI”, and how AI is envisioned to help address social good problems (RQ1). In the Expected Outcomes and Evaluations (Section 4.2), we discuss funding documents characterized outcomes, often communicated through evaluation criteria (RQ2). In the Eligibility (Section 4.3) and Support from Funders (Section 4.4) sections, we shed light on how funding documents described “AI4SG” can be achieved and by whom (RQ3). Table LABEL:tab:funding_doc_mr in Appendix A provides a detailed overview of our dataset, including funding amount, duration, targeted regions, funder types, and target regions.
@@ 4.4.2. More balanced
Tailoring activities and resources to the needs of applicants beyond technical support. Some other funding programs provided broader forms of support, beyond just technical resources and training. For example, as part of the Wellcome Mental Health Data Prize (C8 and C9), Wellcome Trust offered a web-based platform to facilitate knowledge-sharing across teams. The funder also facilitated matchmaking events where “individuals with different knowledge, skills and backgrounds will be introduced to collaborate and develop a proposal together” to enable interdisciplinary collaborations. In another funding call (C7), the Mozilla Foundation described a broad array of support that funded teams would receive: “While the specific types of support are tailored to the unique needs of the awardees and cohorts, providing opportunities for peer learning and community, mentorship, and communications support are key elements of our accompaniment approach.” However, none of the funding documents in our corpus explicitly mentioned training resources and support for local community engagement and participatory design.
A few funding documents provided references for guidelines and academic papers in responsible AI for applicants. These references could serve as a guide for applicants who were not previously familiar with these resources. For example, the funding call for AI innovations to improve Maternal, Sexual, and Reproductive Health and Rights in Sub-Saharan Africa (C10) provided a list of references on responsible AI as resources for applications. This list included Gardner et al. (2022)’s work on the funding processes for responsible AI, which calls for the inclusion of a Trustworthy AI statement section in grant application forms. The application form included a section for responsible AI development, requiring the applicants to elaborate on “how gender equality and inclusion considerations will be integrated into the project design, implementation, monitoring and results,” as well as considerations around “responsible AI, the collection, use and storage of data particularly individually identifiable data” (C10).
@@ 5. Discussion
Our findings extend prior literature on document analysis of AI imaginaries (Zhong et al., 2025), by showing that some AI4SG funding documents similarly reproduce many of the same techno-centric assumptions found in ethics toolkits (Wong et al., 2023b), national AI policies (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022), and military funding solicitations (Widder et al., 2024), but with distinct implications for social-impact work. Like public texts that portray AI as an “autonomous and inevitable force” (Sophie Bennani-Taylor, 2024; Bareis and Katzenbach, 2022; Widder et al., 2024), some AI4SG funding calls in our corpus similarly positioned AI as a transformative solution to large-scale social problems (with limited contextual grounding) and promoted its urgent dissemination to resource-constrained social contexts (like the nonprofit sector and international development) (Section 4.1.1). Some funding documents also mirrored AI ethics toolkits that “frame the work of AI ethics as technical work for individual technical practitioners” (Wong et al., 2023b), providing little to no guidance on how to involve community stakeholders despite encouraging interdisciplinary collaborations. Yet AI4SG funding documents differ from these domains in their explicit aim to advance social good, making the dissonances between expected outcomes and some techno-centric rhetoric particularly stark. We discuss these dissonances in Section 5.1.
At the same time, a subset of AI4SG documents in our dataset adopted more balanced framings, such as situating AI’s potential within contextual understanding of scoped social problems (Section 4.1.2) and requiring domain expertise or contextual knowledge in core project teams (Section 4.3.2). These examples reveal the presence of funding rhetoric that does encourage community-grounded approaches, a departure from dominant AI imaginaries. While there is complexity in how what is written in funding documents influences what actually gets funded, funding rhetoric influences downstream applicants’ behaviors (Peng et al., 2024) and research activities (Widder et al., 2024; Smith et al., 2023). This makes intervening in the design process of funding documents crucial for influencing downstream project approaches before projects are even selected. Building on the more balanced approaches and existing gaps identified in our dataset, we propose opportunities for designing future funding calls for funders, including supporting relationship building, maintenance post-project deployment, and community engagement training for project teams (Section 5.2). Finally, we outline an agenda for meaningful community engagement for funders, and specify the roles that the HCI researchers and practitioners can and should play in helping shape AI4SG funding agendas (Section 5.3)
@@
5.1. Revealing Dissonances Between Expected Outcomes and More Techno-centric Approaches in AI4SG Funding Documents
@@
Many AI4SG funding documents in our dataset expected feasible and sustainable deployment of AI technologies or novel research advancements (Section 4.2). However, more techno-centric approaches promoted in some funding documents may influence downstream approaches taken by AI4SG project teams, impacting AI4SG’s goal of achieving positive and long-lasting change in real-world contexts (Perera, 2024; Tudor et al., 2024). We highlight three prominent aspects below, in conversation with HCI literature on “tech for good” projects.
@@
First of all, funding documents motivating AI’s use for addressing large-scale social problems without instantiating concrete contextual benefits (Section 4.1.1) could lead to premature emphasis on scalability, risking putting such a strong emphasis on creating something that works everywhere that it ends up not working particularly well anywhere. As Hanna and Park note, “scale thinking” necessitates homogenizing users into predefined categories, disregarding the diversity of local contexts (Hanna and Park, 2020). The urgency to scale could incentivize rushed implementations that are poorly suited to diverse contexts (Heeks, 2003, 2002; Dada, 2006). As Burrell and Toyama (2009) put it, “in an effort to solve problems quickly on large scale, a concern for local values, preferences, and opinions may become secondary.” We do not argue that scalable and generalizable approaches should never be pursued, but we suggest that they should be pursued inductively. That is, AI4SG initiatives should focus early efforts on solving specific problems of specific communities effectively and sustainably, and then look for commonalities that could inform the development of generalizable technical capabilities and infrastructures.
@@
Secondly, funding rhetoric that led with assumptions of AI’s deterministic benefits without acknowledging its risks (Section 4.1.1) could reinforce the wider hype and “enchanted determinism” surrounding AI (Campolo and Crawford, 2020). This narrative fosters unrealistic expectations (Lin et al., 2024; Taylor et al., 2013), promotes techno-solutionism, and diverts attention from underlying social issues (Walsham, 2017; Karusala and Kumar, 2017). Furthermore, when funding documents framed AI as a force for good in the Global South or nonprofit sectors deterministically (Section 4.1.1), it risks promoting top-down approaches that cast these groups as lacking and in need of catching up (Erete et al., 2023; Valencia, 1997; Dearden and Tucker, 2016). This framing ignores the unique needs (and assets) of public sectors and the Global South (Hanna and Park, 2020). While corporate donations of computing resources and support (Section 4.4.1) can expand access, they may also reinforce Big Tech’s hegemonic narrative (Collins, 2000; Lukacz, 2024; Nost and Colven, 2022; Seger et al., 2023), addressing surface-level gaps rather than deeper structural issues (Ho et al., 2009b; Burrell and Toyama, 2009; Dearden and Tucker, 2016). Notably, all the funders in our dataset are headquartered in the Global North, even though many funding programs were aimed at applicants or projects located in the Global South. This geographic asymmetry reveals an uneven distribution of hegemonic power (Collins, 2000), as Global North institutions retain disproportionate authority to set funding agendas. These initiatives risk reproducing patterns of saviorism, positioning the Global South as a testing ground for foreign-owned systems while extracting valuable data (Damilare Dosunmu, 2025). These projects also overlook the uneven geography of the AI supply chain; the Global South is already part of AI, but primarily through the underpaid labor that makes AI systems work (Posada, 2022).